OpenStudy (anonymous):

f (x) is the vertical distance in meters traveled by a test rocket in x seconds. determine the slope f represents... what are they asking?? help please.

OpenStudy (phi):

is this calculus or geometry ? is there more background on the question you left out ?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

that's all there is. word for word

OpenStudy (anonymous):

The slope f represents velocity. f(x): distance as a function of time, i.e. distance / time = velocity.

OpenStudy (phi):

it is ambiguous to me... but if we say y= f(x) then you could say the slope of the curve at point x is f'(x) = dy/dx

OpenStudy (anonymous):

To put it simply: any graph is a function = f(x) [f = function, of (x)] of one variable. To say, y is a function of x. Meaning the values of y (dependent variable) are derived from the values of x (independent variable). "f (x) is the vertical distance in meters traveled by a test rocket in x seconds. " This statement assumes that f(x), which is vertical distance, is DEPENDENT on time (x seconds). Hence, you would have the y-axis as distance, and x-axis as time. Assuming a line bounded by these two axis, and the gradient / slope of this line is determined by taking dy (change in y) / dx (change in x). Which goes to say: y / x = distance / time = velocity.